Two trains leave simultaneously from stations that are 360 miles apart. If one train is traveling at 40 miles per hour and the other is traveling 50 miles per hour, how far (along the track) from the nearest station will they meet?
Found 2 solutions by oberobic, nerdybill:
Answer by oberobic(2304)

I assume the trains are heading toward each other from the two different stations.
That means they are converging at a combined speed of 90 mph.
When they meet, they will have covered the 360 miles.
The fundamental distance equation is d = rt.
In this case we know d = 360 miles and r = 40+50 = 90 mph. Therefore,
t = 360/90 = 4 hr
The train going 40 mph will have covered 4*40 = 160 miles at the time they meet.
The train going 50 mph will have traveled 4*50 = 200 miles at the time they meet.
160+200 = 360 miles, so that checks.
Re-reading the question, it asks for how far from the nearest station they will be.
Well, both of them will be 160 miles from the nearest station. One train will heading toward the nearest station and the other will be heading away from it.
Done.

Answer by nerdybill(7384)

wo trains leave simultaneously from stations that are 360 miles apart. If one train is traveling at 40 miles per hour and the other is traveling 50 miles per hour, how far (along the track) from the nearest station will they meet?
.
apply the distance formula: d=rt
.
first, determine the time traveled before they mee50t:
Let x = time (hours) before they meet
then
40x + 50x = 360
90x = 360
x = 4 hours
.
distance from nearest station is equivalent to distance traveled by the slower train:
distance = 40(4) = 160 miles